rssLink RSS for all categories
 
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_green
icon_green
icon_red
icon_red
icon_red
icon_green
icon_green
icon_green
icon_blue
icon_orange
icon_red
icon_green
icon_blue
icon_red
icon_orange
icon_green
icon_red
icon_red
icon_blue
icon_orange
icon_green
icon_green
icon_red
icon_green
icon_blue
icon_green
icon_green
 

FS#4460 — Amsterdam <> Varsovie / Francfort <> Varsovie

Attached to Project— Reseau Internet et Baies
Incident
Tout le réseau
CLOSED
100%
Le routeur sur Varsovie ne répond plus. Nous investiguons.
Date:  Wednesday, 11 August 2010, 10:09AM
Reason for closing:  Done
Additional comments about closing:  Le trafic est rétabli.
Comment by OVH - Tuesday, 10 August 2010, 18:29PM

Il s'agit d'un problème sur les liaisons Amsterdam<>Varsovie et Francfort<>Varsovie. Nous redirigons le trafic plix vers d'autres peerings/transit afin d'éviter la saturation des liens de backups Francfort<>Prague et Prague<>Varsovie.


Comment by OVH - Tuesday, 10 August 2010, 19:12PM

une fibre coupée en DE et il ne faudrait pas
espérer de rétablissement avant minuit.

Fin du mois, on va basculer le lien Frankfurt/Varsovie par Prague au lieu par Berlin.


Comment by OVH - Tuesday, 10 August 2010, 21:45PM

Update 2 (18:20 UTC/ 20:20 CEST): Running outage time 2h15min
The engineers have arrived on break location which is within a 2 km distance from the city of Braunschweig/ Germany.


Comment by OVH - Tuesday, 10 August 2010, 21:56PM

Update 3 (19:15 UTC/ 21:15 CEST): Running outage time 3h10min
Correction from previous update- the engineers had arrived at the location for OTDR measurement. The cut itself has been located 2,7 km from Braunschweig. In the area there are ongoing construction works on a motorway. Engineers are now heading for the break location. No ETA.


Comment by OVH - Tuesday, 10 August 2010, 23:37PM

2 des 4 circuits impactés ont été reroutés sur un autre cable. Nous avons récupéré 1x10G Varsovie<>Amsterdam et 1x10G Varsovie<>Francfort


Comment by OVH - Wednesday, 11 August 2010, 01:55AM

Update 6 (23:20 UTC/ 01:20 CEST): Running outage time 7h15min
The solution of this fiber cut willbe optical bypass. Engineers are still digging the cable (one end completed, works on second point still ongoing) First ETR is 3-4h from now.